# A Pipeline to Assess Merging Methods via Behavior and Internals

### v1

@@ 1 Introduction
With the rise of competitive open-weight models (Dubey et al., 2024; Jiang et al., 2024), adapting large language models (LLMs) to specific use cases has become common practice. One approach, fine-tuning, enables adaptation to particular domains, such as mathematics or coding Lewkowycz et al. (2022). Still, it suffers from high computational costs and the risk of catastrophic forgetting (Luo et al., 2023). When a model must handle a variety of tasks, fine-tuning becomes more complex Lu et al. (2024). Model merging offers methods that overcome some of these issues. These methods combine multiple models into a single one, thereby enhancing performance across the tasks for which the individual parent models were fine-tuned (Yang et al., 2024).
Existing work in model merging primarily focuses on techniques to improve the performance of merged models. For example, Lu et al. (2024) explored strategies for merging domain-adapted parent models, such as those in materials science and engineering, to efficiently transfer this domain adaptation to other models. In addition, Dziadzio et al. (2024) introduces the concept of temporal model merging, addressing the challenge of integrating knowledge from multiple parent models trained on various tasks over time. Furthermore, Goddard et al. (2025) introduces MergeKit, a merging toolkit supporting various methods.
@@
To answer this, we present an evaluation pipeline (section 3) with three stages: 1) we merge multiple parent models with different strategies, 2) we evaluate both the behavior and internals of the parent and resulting merged models, and 3) we connect the findings from both evaluations. We showcase this pipeline through experiments on the Qwen2.5 model family (Qwen et al., 2025), which merges instruction fine-tuned and domain-adapted models. Our results show that while merged models generally perform between the parent models, their encoded linguistic competence—particularly for morphology and syntax—can be stronger than the parent ones’. At the same time, weak rank correlations between behavior and internals underscore the need for comprehensive evaluations, as single perspectives are insufficient to capture the full complexity of language models (Hu and Levy, 2023; Waldis et al., 2025), for and beyond model merging.
@@ 2 Related works
Model merging enables the efficient combination of specialized models (Yang et al., 2024). This includes simpler methods like linearly averaging multiple parent models Wortsman et al. (2022) or spherical interpolation (SLERP) [^2] https://github.com/Digitous/LLM-SLERP-Merge. More sophisticated methods try to locate task-specific regions to better preserve specific skills of the parent models, popular examples are Task Arithmetic (Ilharco et al., 2023), Ties (Yadav et al., 2023), or Dare-Ties (Yu et al., 2024). The general popularity of such methods is evident in the development of toolboxes that easily merge two or more parent models, such as MergeKit (Goddard et al., 2024), resulting in thousands of merged models available on Huggingface.
@@
A predominant line of research focuses on evaluating language models based on their behavior. This includes general language understanding (GLUE; Wang et al. 2018 or SuperGLUE; Wang et al. 2019), question answering (SQuAD; Rajpurkar et al. 2016), or more broad evaluations, as done in HELM (Liang et al., 2023) or evaluation harness (Gao et al., 2024). Moreover, research also focused on comprehensively assessing specific domains, like factuality (Chen et al., 2023; Muhlgay et al., 2024), medical texts (Bedi et al., 2025), or legal reasoning (Guha et al., 2023). While these benchmarks focus on model behavior, there has been little comprehensive work addressing model internals. Conneau et al. (2018) introduce a benchmark with ten tasks to assess linguistic properties of sentence representations, Warstadt et al. (2020) presents a collection of minimal pairs to examine whether language models can internally differentiate between linguistic acceptable and unacceptable sentences, and Waldis et al. (2024) introduced, with Holmes, a benchmark that studies the linguistic competence of language models based on their internals across five linguistic phenomena and more than 160 distinct probing tasks.
@@ 3.1 Merging methods
In a first step, this pipeline uses MergeKit (Goddard et al., 2024) to combine two parent models. Table 1 provides an overview of these methods, which vary in complexity, and compares them based on a few essential features. Note that performance always depends on the configuration and model, such that Linear can work better than DARE TIES under certain conditions.
@@
averages corresponding weights of multiple models, often fine-tuned from a common base with different hyperparameters. This method stands out because of its low complexity and very low difficulty when implemented with Mergekit, making it highly accessible to practitioners. The approach maintains low power usage during both inference and merging processes, while supporting multi-model combinations effectively. Key Application/Strength are simplicity, improved accuracy, and robustness without an increase in inference costs for similar models. Despite its simplicity, performance can vary significantly depending on configuration and, in some cases, may be limited compared to more complex methods Wortsman et al. (2022).
@@
interpolates weights between two models along spherical paths, with the aim of a smoother blend than Linear. This method operates with moderate complexity while maintaining low power consumption during both inference and merging. However, it is limited to only two models in practice, making it less suitable for multi-model scenarios. The implementation difficulty of Mergekit remains low, making it accessible for most applications. Key Application/Strength are effective two-model merging, potentially finding lower loss barrier paths. Performance typically exceeds linear methods when properly configured, offering a balanced approach between simplicity and effectiveness Shoemake (1985); Chris Alexiuk (2024).
@@
computes "task vectors" (fine-tuned base weights) and adds them to a base model to combine capabilities. This approach maintains moderate complexity while effectively supporting multi-model merging scenarios. The method benefits from low power usage during both inference and merging phases, with relatively low implementation difficulty in Mergekit frameworks. Key Application/Strength are flexible multitasking combination in model editing. Performance generally surpasses Linear approaches, providing a practical solution to combine diverse model capabilities without significant computational overhead Ilharco et al. (2023).
@@
trims insignificant parameter changes, selects a dominant sign for conflicts, and merges only aligned parameters. This method introduces high complexity in its implementation, which requires careful configuration and understanding of the interaction of the parameters. Despite the increased complexity, it maintains low power usage during inference and merging, while supporting multi-model combinations. Key Application/Strength are the reduction of parameter interference, especially for sign conflicts, in multi-model merging. Performance typically exceeds SLERP methods when properly implemented, making the additional complexity worthwhile for demanding applications Yadav et al. (2023).
@@
uses the DARE (Drop And REscale) scheme to sparsify task vectors (randomly drops and rescales parameters) before applying TIES-merging. This advanced method operates with high complexity, building on the TIES framework with additional parameter sparsification techniques that require expertise in both DARE and TIES methodologies. Power consumption remains low during both inference and merging, while supporting comprehensive multi-model integration. Key Application/Strength are the further mitigated interference by reducing parameter density before TIES. Performance generally surpasses SLERP methods and can exceed TIES in many configurations, representing a state-of-the-art approach to model merging Yadav et al. (2023); Yu et al. (2024).
@@ 3.2 Evaluation
In a second step, our pipeline automatically evaluates both the parent models and the resulting merged ones. This includes a behavioral evaluation using the eval-harness library (Gao et al., 2024) and the streamlined and efficient version of Holmes (Waldis et al., 2024), which comprehensively assesses information about linguistic phenomena within model internals.
@@ 3.2.1 LM-Harness-Evaluation
The LM-Harness-Evaluation is a comprehensive evaluation framework to assess language models (Gao et al., 2024). It includes a variety of tasks in favor of a standardized and reproducible evaluation of these models. For our study, we use the following tasks also used in the OpenLLM evaluation leaderboard (Fourrier et al., 2024):
@@ 3.2.2 Holmes-Evaluation
Using Holmes, we evaluate the internal representations using classifier based probing (Belinkov, 2022) Specifically, we use Flash-Holmes, a streamlined and efficient version of the benchmark that preserves the effectiveness of the evaluation, even substantially reducing the number of instances. It assesses a range of linguistic phenomena, including morphology, syntax, semantics, reasoning, and discourse, using more than 160 unique probing tasks. This involves training simple linear classifiers on the internal representations of the models’ last layer to predict specific linguistic properties, thereby assessing how well the model encodes this information Hewitt and Liang (2019); Voita and Titov (2020); Waldis et al. (2024).
@@ 4 Evaluation Results
We merge Instruct with Coder and Math models from the 7B Qwen-2.5 family Qwen et al. (2025). Leveraging both the LM-Harness and Holmes frameworks reveals a significant and consistent divergence in how model merging impacts external behavior versus internal representations. While downstream tasks, the best merged models perform between the two parent models, the inherent linguistic competence encoded within the model internals tends to increase, particularly when merging Instruct and Math.
@@ 4.1 Model Behavioral Evaluation
These results consistently demonstrate that merged models can not match the performance of one of their parent models, but mostly perform between them. As an exception, in both Instruct + Math and Instruct + Coder experiments, the merged models fail to reach at least one parent model’s performance for MATH Hard. These results align with the general low performance of domain-adapted models and previous results, which underscore the importance of instruction-tuning in combination with domain adaptation (Beeching et al., 2023).
@@ 4.2 Model Internal Evaluation
In the next step, we compare the different merging methods regarding the information encoded in the resulting models. Similarly, to the behavioral evaluation, we find that simpler methods (Liner or SLERP) generally preserve more information than more complex methods. This effect is not as pronounced as when evaluating model behavior. Notably, we find the most difference among models (parent and merged ones) in more formal phenomena, like syntax and morphology. However, there are fewer differences in phenomena that are intuitively linked to actively using language (reasoning, semantics, or discourse), which we often attribute to language models. This improvement for formal phenomena also suggests that merging can effectively transfer the information gain from seeing more tokens in further pre-training steps, as shown in (Waldis et al., 2024).
@@ 6 Conclusion
Given the widespread availability of evaluation resources, we limited this study to the English language. Thereby, the presented pipeline is not directly applicable to multilingual merging methods (Tao et al., 2024).

### v2

@@ 1 Introduction
With the rise of competitive open-weight models (Dubey2024TheL3; jiang2024mixtral), adapting large language models (LLMs) to specific use cases has become common practice. One approach, fine-tuning, enables adaptation to particular domains, such as mathematics or coding lewkowycz2022solvingquantitativereasoningproblems. Still, it suffers from high computational costs and the risk of catastrophic forgetting (DBLP:journals/corr/abs-2308-08747). When a model must handle a variety of tasks, fine-tuning becomes more complex lu2024finetuninglargelanguagemodels. Model merging offers methods that overcome some of these issues. These methods combine multiple models into a single one, thereby enhancing performance across the tasks for which the individual parent models were fine-tuned (yang2024modelmergingllmsmllms).
Existing work in model merging primarily focuses on techniques to improve the performance of merged models. For example, lu2024finetuninglargelanguagemodels explored strategies for merging domain-adapted parent models, such as those in materials science and engineering, to efficiently transfer this domain adaptation to other models. In addition, dziadzio2024mergemultimodalmodelstime introduces the concept of temporal model merging, addressing the challenge of integrating knowledge from multiple parent models trained on various tasks over time. Furthermore, goddard2025arceesmergekittoolkitmerging introduces MergeKit, a merging toolkit supporting various methods.
@@
To answer this, we present an evaluation pipeline (section 3) with three stages: 1) we merge multiple parent models with different strategies, 2) we evaluate both the behavior and internals of the parent and resulting merged models, and 3) we connect the findings from both evaluations. We showcase this pipeline through experiments on the Qwen2.5 model family (qwen2025qwen25technicalreport), which merges instruction fine-tuned and domain-adapted models. Our results show that while merged models generally perform between the parent models, their encoded linguistic competence—particularly for morphology and syntax—can be stronger than the parent ones’. At the same time, weak rank correlations between behavior and internals underscore the need for comprehensive evaluations, as single perspectives are insufficient to capture the full complexity of language models (hu-levy-2023-prompting; waldis2025aligned), for and beyond model merging.
@@ 2 Related works
Model merging enables the efficient combination of specialized models (yang2024modelmergingllmsmllms). This includes simpler methods like linearly averaging multiple parent models wortsman2022modelsoupsaveragingweights or spherical interpolation (SLERP) [^2] https://github.com/Digitous/LLM-SLERP-Merge. More sophisticated methods try to locate task-specific regions to better preserve specific skills of the parent models, popular examples are Task Arithmetic (ilharco2023editingmodelstaskarithmetic), Ties (yadav2023tiesmergingresolvinginterferencemerging), or Dare-Ties (yu2024languagemodelssupermario). The general popularity of such methods is evident in the development of toolboxes that easily merge two or more parent models, such as MergeKit (goddard-etal-2024-arcees), resulting in thousands of merged models available on Huggingface.
@@
A predominant line of research focuses on evaluating language models based on their behavior. This includes general language understanding (GLUE; wang-etal-2018-glue or SuperGLUE; 10.5555/3454287.3454581), question answering (SQuAD; rajpurkar-etal-2016-squad), or more broad evaluations, as done in HELM (liang2022holistic) or evaluation harness (eval-harness). Moreover, research also focused on comprehensively assessing specific domains, like factuality (chen2023felm; DBLP:conf/eacl/MuhlgayRMLRBALSS24), medical texts (bedi2025medhelm), or legal reasoning (guha2023legalbench). While these benchmarks focus on model behavior, there has been little comprehensive work addressing model internals. conneau-etal-2018-cram introduce a benchmark with ten tasks to assess linguistic properties of sentence representations, warstadt-etal-2020-blimp-benchmark presents a collection of minimal pairs to examine whether language models can internally differentiate between linguistic acceptable and unacceptable sentences, and waldis2024holmesbenchmarkassesslinguistic introduced, with Holmes, a benchmark that studies the linguistic competence of language models based on their internals across five linguistic phenomena and more than 160 distinct probing tasks.
@@ 3.1 Merging methods
In a first step, this pipeline uses MergeKit (goddard-etal-2024-arcees) to combine two parent models. Table 1 provides an overview of these methods, which vary in complexity, and compares them based on a few essential features. Note that performance always depends on the configuration and model, such that Linear can work better than DARE TIES under certain conditions.
@@
averages corresponding weights of multiple models, often fine-tuned from a common base with different hyperparameters. This method stands out because of its low complexity and very low difficulty when implemented with Mergekit, making it highly accessible to practitioners. The approach maintains low power usage during both inference and merging processes, while supporting multi-model combinations effectively. Key Application/Strength are simplicity, improved accuracy, and robustness without an increase in inference costs for similar models. Despite its simplicity, performance can vary significantly depending on configuration and, in some cases, may be limited compared to more complex methods wortsman2022modelsoupsaveragingweights.
@@
interpolates weights between two models along spherical paths, with the aim of a smoother blend than Linear. This method operates with moderate complexity while maintaining low power consumption during both inference and merging. However, it is limited to only two models in practice, making it less suitable for multi-model scenarios. The implementation difficulty of Mergekit remains low, making it accessible for most applications. Key Application/Strength are effective two-model merging, potentially finding lower loss barrier paths. Performance typically exceeds linear methods when properly configured, offering a balanced approach between simplicity and effectiveness Animating_rotation_with_quaternion_curves; NVIDIATechnicalBlog_2024.
@@
computes "task vectors" (fine-tuned base weights) and adds them to a base model to combine capabilities. This approach maintains moderate complexity while effectively supporting multi-model merging scenarios. The method benefits from low power usage during both inference and merging phases, with relatively low implementation difficulty in Mergekit frameworks. Key Application/Strength are flexible multitasking combination in model editing. Performance generally surpasses Linear approaches, providing a practical solution to combine diverse model capabilities without significant computational overhead ilharco2023editingmodelstaskarithmetic.
@@
trims insignificant parameter changes, selects a dominant sign for conflicts, and merges only aligned parameters. This method introduces high complexity in its implementation, which requires careful configuration and understanding of the interaction of the parameters. Despite the increased complexity, it maintains low power usage during inference and merging, while supporting multi-model combinations. Key Application/Strength are the reduction of parameter interference, especially for sign conflicts, in multi-model merging. Performance typically exceeds SLERP methods when properly implemented, making the additional complexity worthwhile for demanding applications yadav2023tiesmergingresolvinginterferencemerging.
@@
uses the DARE (Drop And REscale) scheme to sparsify task vectors (randomly drops and rescales parameters) before applying TIES-merging. This advanced method operates with high complexity, building on the TIES framework with additional parameter sparsification techniques that require expertise in both DARE and TIES methodologies. Power consumption remains low during both inference and merging, while supporting comprehensive multi-model integration. Key Application/Strength are the further mitigated interference by reducing parameter density before TIES. Performance generally surpasses SLERP methods and can exceed TIES in many configurations, representing a state-of-the-art approach to model merging yadav2023tiesmergingresolvinginterferencemerging; yu2024languagemodelssupermario.
@@ 3.2 Evaluation
In a second step, our pipeline automatically evaluates both the parent models and the resulting merged ones. This includes a behavioral evaluation using the eval-harness library (eval-harness) and the streamlined and efficient version of Holmes (waldis2024holmesbenchmarkassesslinguistic), which comprehensively assesses information about linguistic phenomena within model internals.
@@ 3.2.1 LM-Harness-Evaluation
The LM-Harness-Evaluation is a comprehensive evaluation framework to assess language models (eval-harness). It includes a variety of tasks in favor of a standardized and reproducible evaluation of these models. For our study, we use the following tasks also used in the OpenLLM evaluation leaderboard (open-llm-leaderboard-v2):
@@ 3.2.2 Holmes-Evaluation
Using Holmes, we evaluate the internal representations using classifier based probing (belinkov-2022-probing) Specifically, we use Flash-Holmes, a streamlined and efficient version of the benchmark that preserves the effectiveness of the evaluation, even substantially reducing the number of instances. It assesses a range of linguistic phenomena, including morphology, syntax, semantics, reasoning, and discourse, using more than 160 unique probing tasks. This involves training simple linear classifiers on the internal representations of the models’ last layer to predict specific linguistic properties, thereby assessing how well the model encodes this information hewitt2019designinginterpretingprobescontrol; voita2020informationtheoreticprobingminimumdescription; waldis2024holmesbenchmarkassesslinguistic.
@@ 4 Evaluation Results
We merge Instruct with Coder and Math models from the 7B Qwen-2.5 family qwen2025qwen25technicalreport. Leveraging both the LM-Harness and Holmes frameworks reveals a significant and consistent divergence in how model merging impacts external behavior versus internal representations. While downstream tasks, the best merged models perform between the two parent models, the inherent linguistic competence encoded within the model internals tends to increase, particularly when merging Instruct and Math.
@@ 4.1 Model Behavioral Evaluation
These results consistently demonstrate that merged models can not match the performance of one of their parent models, but mostly perform between them. As an exception, in both Instruct + Math and Instruct + Coder experiments, the merged models fail to reach at least one parent model’s performance for MATH Hard. These results align with the general low performance of domain-adapted models and previous results, which underscore the importance of instruction-tuning in combination with domain adaptation (Pham_OpenLLM_Operating_LLMs_2023).
@@ 4.2 Model Internal Evaluation
In the next step, we compare the different merging methods regarding the information encoded in the resulting models. Similarly, to the behavioral evaluation, we find that simpler methods (Liner or SLERP) generally preserve more information than more complex methods. This effect is not as pronounced as when evaluating model behavior. Notably, we find the most difference among models (parent and merged ones) in more formal phenomena, like syntax and morphology. However, there are fewer differences in phenomena that are intuitively linked to actively using language (reasoning, semantics, or discourse), which we often attribute to language models. This improvement for formal phenomena also suggests that merging can effectively transfer the information gain from seeing more tokens in further pre-training steps, as shown in (waldis2024holmesbenchmarkassesslinguistic).
@@ 6 Conclusion
Given the widespread availability of evaluation resources, we limited this study to the English language. Thereby, the presented pipeline is not directly applicable to multilingual merging methods (tao-etal-2024-unlocking).